# Edinburgh 2025/29

### v1

@@ 5 Conclusions
We have presented a determination of higher twist corrections in DIS, and linear power corrections in single-inclusive jets and dijets in a joint fit with global PDFs. We find good agreement with earlier results for higher twist. This is the first time a determination of power corrections in jets has been performed in the context of PDF fits, but we find agreement with theoretical expectations. More specifically, while the higher twist corrections are mainly taken care of by cuts in $Q^{2}$ of a few GeV2, and in $W^{2}$ at a few tens of GeV2, we find that the effect of the linear power corrections in jets can extend up to hard scales of several hundred GeV. This means in particular that when incorporating LHC data on single-inclusive jets in global PDF fits, it may be necessary for PDFs accurate to within 1% to either cut some of the low $p_{T}$ data, or include a linear power correction in the theoretical predictions (or indeed both, similar to the strategy for DIS).
@@
Since the effects of higher twist and linear power corrections in jets can be relevant at the current level of precision for various processes, they should be accounted for in future global PDF determinations that hope to achieve a precision of 1% in LHC observables. The empirical determination of power corrections in Drell-Yan and vector boson production processes, and the possible linear power corrections in top production [Makarov:2023uet], are left for future studies.
The PDF sets with power corrections are made available in LHAPDF format [Buckley:2014ana] on the NNPDF website
@@
After this work was completed, two other studies of higher twist in the context of PDF fits and $\alpha_{s}$ appeared [Harland-Lang:2025wvm, Alekhin:2025qdj]. Neither of these consider higher twist for charge current processes, or linear power corrections to jet processes.
@@ Acknowledgements:
We would like to thank Gavin Salam for discussions at Gargnano in September 2023, where he emphasised the possible significance of linear power corrections to jet observables in global PDF fits. We would also like to thank the members of the NNPDF collaboration for insightful discussions during the course of this work. Finally, we are grateful to Felix Hekhorn for providing valuable feedback on the first draft of this manuscript. R.D.B. and R.S. are supported by the Science and Technology Facilities Council (STFC) via grant awards ST/T000600/1 and ST/X000494/1.

### v2

@@ 5 Conclusions
We have presented a determination of higher twist corrections in DIS, and linear power corrections in single-inclusive jets and dijets in a joint fit with global PDFs. We find good agreement with earlier results for higher twist. This is the first time an empirical determination of power corrections in jets has been performed in the context of PDF fits, but we find agreement with theoretical expectations, and in particular with estimates of hadronization corrections made using final state Monte Carlos.
While the higher twist corrections are mainly taken care of by cuts in $Q^{2}$ of a few GeV2, and in $W^{2}$ at a few tens of GeV2, we find that the effect of the linear power corrections in jets can extend up to hard scales of several hundred GeV. This means in particular that when incorporating LHC data on single-inclusive jets in global PDF fits, it may be necessary for PDFs accurate to within 1% to either cut some of the low $p_{T}$ data, or include a linear power correction in the theoretical predictions (or indeed both, similar to the strategy for DIS). The size of the power correction could be reduced by first correcting the data (from hadron to parton level) using a correction factor estimate using final state Monte Carlos, if such factors were determined reliably and consistently across all available jet datasets.
@@
Since the effects of higher twist and linear power corrections in jets can be relevant at the current level of precision for various processes, they should be accounted for in future global PDF determinations that hope to achieve a precision of 1% in LHC observables. The empirical determination of power corrections in Drell-Yan and vector boson production processes, and the possible linear power corrections in top production [65], are left for future studies.
The PDF sets with power corrections are made available in LHAPDF format [45] on the NNPDF website
@@ Acknowledgements:
We would like to thank Gavin Salam for discussions at Gargnano in September 2023, where he emphasised the possible significance of linear power corrections to jet observables in global PDF fits. We would also like to thank the members of the NNPDF collaboration for insightful discussions during the course of this work. Finally, we are grateful to Felix Hekhorn and Emanuele Nocera for providing valuable feedback on the earlier drafts of this manuscript. R.D.B. and R.S. are supported by the Science and Technology Facilities Council (STFC) via grant awards ST/T000600/1 and ST/X000494/1.